Find Torison r=(a (3u, u³), 3au², 3a (3u+u³)



1
Expert's answer
2021-11-16T15:32:59-0500
"r=\langle a (3a-u^3), 3au^2, 3a (3u+u^3)\rangle"

"r'=\langle -3au^2, 6au, 9a (1+u^2)\rangle"

"r''=\langle -6au, 6a, 18au\rangle"

"r'''=\langle -6a, 0, 18a\rangle"

"r'\times r''=\begin{vmatrix}
 i & j & k \\
 -3au^2 & 6au & 9a(1+u^2)\\
-6au & 6a & 18au
\end{vmatrix}"

"=i(108a^2u^2-54a^2(1+u^2))"

"-j(-54a^2u^3+54a^2u(1+u^2))"

"+k(-18a^2u^2+36a^2u^2)"

"=54a^2(u^2-1)i-54a^2uj+18a^2u^2k"

"(r'\times r'')\cdot r'''=54a^2(u^2-1)(-6a)-0+18a^2u^2(18a)"

"=-324a^3u^2+324a^3+324a^3u^2=324a^3"

"||r'\times r''||^2=(54a^2(u^2-1))^2+(-54a^2u)^2+(18a^2u^2)^2"

"=324a^4(9u^4-18u^2+9+9u^2+u^4)"

"=324a^4(10u^4-9u^2+9)"

"\tau=\dfrac{(r'\times r'')\cdot r'''}{||r'\times r''||^2}"

"=\dfrac{324a^3}{324a^4(10u^4-9u^2+9)}=\dfrac{1}{a(10u^4-9u^2+9)}"

"\tau=\dfrac{1}{a(10u^4-9u^2+9)}"